OpenStudy (anonymous):

Evaluate: [(51 + 3) − 32] ÷ 9 • 2 52 13.3 2.5 10

OpenStudy (anonymous):

Go solving the brackets in given order [(51 + 3) − 32] ÷ 9 • 2 = [54 − 32] ÷ 9 • 2 = [22] ÷ 9 • 2 = 22 divided by 9 then multiplied by 2 = 44 divided by 9 = 44/9 In decimal format, it is 4.88888888 ......... you can round it as 4.889, 4.89. 4.9 approx.

OpenStudy (anonymous):

.... so which one? that confushed me

OpenStudy (anonymous):

not sure which one since i did the order of operations should be 4.9 for order of operations you use PEMDAS Parentheses Exponents Multiplication Division addition subtraction

OpenStudy (anonymous):

@ninab731 did the problem correctly.... can you check to see if you wrote the problem correctly?

OpenStudy (anonymous):

i got thesame 44/6

OpenStudy (anonymous):

44/9*

OpenStudy (anonymous):

is it written down correctly

OpenStudy (anonymous):

or wrote the correct choices ...

OpenStudy (anonymous):

oh yeah?

OpenStudy (anonymous):

wait why..

OpenStudy (anonymous):

?? none of the answers you have written are correct for that problem

OpenStudy (anonymous):

that is what is given.

OpenStudy (anonymous):

then check with teacher or something dont know do order of operations for the problem (51 + 3) − 32] ÷ 9 • 2 = [54 − 32] ÷ 9 • 2 = [22] ÷ 9 • 2 = 22 divided by 9 then multiplied by 2 = 44 divided by 9 = 44/9 In decimal format, it is 4.88888888 ......... you can round it as 4.889, 4.89. 4.9 approx <---- should be the answer

OpenStudy (anonymous):

its supposed to be 3^2